# Instability and stress fluctuations of a probe driven through a worm-like micellar fluid

### v1

@@ Abstract
A particle moving through a worm-like micellar fluid (WLM) shows instability and large fluctuations beyond a threshold. Despite many detailed studies, a direct measurement of the time-dependent stress on the probe particle remains unexplored. To address this, we have designed a measuring geometry coupled with a commercial rheometer to study the dynamics of a cylindrical probe through a WLM system of 2 wt.% cetyltrimethyl ammonium tosylate(CTAT) + 100 mM sodium chloride(NaCl) for a wide range of velocity and stress scales. We map out the in-situ velocity distribution using particle imaging velocimetry. Beyond a certain velocity threshold, we observe large stress fluctuation events with gradual stress build-up followed by sudden stress drop indicating the storage and release of elastic energy. The length scale constructed from the stress build-up time scale and the probe’s velocity match the length scale of extensile deformation in the sample just before the stress drop, further confirming the strong correlation of such storage and release of energy with the unstable motion of the probe. Interestingly, despite their significant difference in magnitudes, the Weissenberg number () for the onset of flow instability obtained from the shear and extensile components remains almost the same. We also find that the turbulent motion of the probe at higher results from the complex mixing of the stick-slip events originating from the partial release of the stored elastic energy. Further, we show that the magnitude of the stick-slip events depends on the detailed micellar structure and dynamics controlled by salt concentration and temperature.
@@ I Introduction
The motion of a particle moving under a constant applied force through a viscous fluid attains a terminal velocity when the viscous drag completely balances the applied force. Such dynamics leading to terminal velocity are well understood[1]. On the other hand, particle motion through visco-elastic worm-like micellar (WLM) fluids can be quite complex, and terminal velocity may not exist due to various fluid instabilities[2, 3, 4, 5]. Understanding such complex motion in WLM systems is important not only for a wide range of applications in oil drilling, hydraulic fracturing, and manufacturing cosmetic products, but they also present interesting challenges from the fundamental physics perspectives [6, 7, 8, 9].
@@
Worm-like micelles (WLMs) are long polymer-like structures formed by the self-assembly of surfactant molecules in aqueous solution. In the semi-dilute regime, the density of these polymer-like structures becomes high enough to form entanglement between them. Such entangled structures give rise to long-range correlations in the system resulting in visco-elasticity. Despite many similarities between these systems and polymeric liquids, the major difference between these two classes of systems stems from the fact that WLM can break and recombine due to thermal fluctuations giving rise to additional relaxation modes. This results in Maxwellian behaviour in a wide range of WLM systems with a well-defined relaxation time scale [10]. Non-linear flows of WLM systems have been extensively studied using shear rheometry as they exhibit striking flow instabilities, such as shear banding, elastic turbulence, rheochaos at substantially low values of Reynolds numbers but at high Weissenberg numbers [11, 12, 13, 14, 15, 16, 17, 18, 19]. Such instabilities are driven by elasticity rather than inertia. Nonetheless, interesting inertio-elastic flow instabilities have also been reported in these systems when both and becomes comparable [20, 21]. Although these studies provide valuable insights regarding the coupling of microstructure and flow in WLM systems, such rheological studies are inadequate to predict/describe the complex motion of a probe through WLM fluids mostly due to the wide differences in flow geometries.
@@
Earlier experimental studies have reported the unstable motion of a sphere settling under gravity through WLM fluids. These include the study by Jayraman and Belmonte[2] using CTAB-NaSal based WLM fluid. They observe that beyond a certain value of Deborah number 45, the velocity of the sphere shows large fluctuations. They have attributed the origin of such fluctuations to the formation and breakage of shear-induced structures around the sphere. More recent studies by Mohammadigoushki and Muller[22] and Zhang and Muller[23] point out the role of extensional deformation in the wake of the moving sphere giving rise to instability beyond the certain value of the extensional Weissenberg number. Using in-situ flow birefringence studies Chen and Rothstein[3] suggest that the flow instability originates from the flow-induced rupture of stretched micelles at the rear end of the moving sphere. They also directly observe filament rupture using extensional rheometry. The experimental investigation of Wu and Mohammadigoushki[5] further supports a similar mechanism of micellar breakage. Kumar et al.[4] show unstable flows in CTAT + NaCl based WLM system where they report a periodic burst of damped oscillations in the sphere’s velocity. The nature of such oscillations is quite distinct from the similar studies mentioned above. This implies that the microscopic micellar structure plays an important role in determining the nature of instability. Nonetheless, the steep rise in velocity of the settling sphere at the onset of a burst also suggests a sudden structural breakage. Interestingly, a recent bulk rheological study on the CTAT-NaCl WLM system highlights the role of micellar breakage in the elasticity-driven Taylor-Couette instability in these systems[16]. Recent numerical simulation based on two-species VCM model[24] for WLM also supports the micellar rupture in the wake of a moving sphere as the mechanism of instability[25]. The role of surface roughness of the settling sphere has also been investigated in WLM systems[26]. The structure of CTAT-based WLM systems is complex and highly sensitive to the nature and concentration of added salt which in turn affect their rheological response significantly[27, 28]. Very recent studies indicate that WLM systems having almost identical linear rheological properties can display significantly different non-linear dynamical behaviour[29].
@@
Despite these detailed studies, a direct measurement of temporal stress on the probe during the transition from stable to unstable motion by systematically varying the flow rate remains unexplored. Based on pressure drop measurements in an array of cylinders [30] and position fluctuations of a flexible elastomer [31] placed in a flow-field of WLM fluids, the role of stress variation during the flow instabilities have been indicated. Nonetheless, stress measurement with high spatial and temporal resolution is crucial to verify the proposed mechanisms of micellar scission in giving rise to unstable motion through WLM systems. Furthermore, such controlled experiments are required to understand the connection of the instability near the onset point to the chaotic or turbulent motion at much higher values of as reported in the earlier studies for CTAT-NaCl systems.
@@
In this work, we use a home-built setup coupled with a commercial rheometer to study the stress response of a cylindrical probe driven through the WLM system with constant velocity. Our setup not only mimics the motion of small probes settling under gravity in a more controlled manner but also has other significant advantages: For a given WLM system, the wide range of velocity/force scales spanning several decades that can be probed/measured in our pin-driving setup can not be matched in conventional ball-drop experiments. Similarly, in our setup the stress fluctuation data can be acquired, as long as the sample remains stable. This allows us to perform accurate statistical analysis over very long time scales. Thus, our setup can potentially verify the applicability of the existing macroscopic rheological models of flow instability, transient dynamics, shear banding etc.[32, 33, 34] in the context of unstable motion of a small probe through similar systems. This setup also allows us to probe the in-situ local deformation field using PIV and light scattering measurements. We observe the formation of a turbid band in the wake of the probe, just below the onset velocity of the probe corresponding to large stress fluctuations. We find that such a turbid band corresponds to an extended wake and its attachment-detachment dynamics with the probe is strongly correlated with the nature of the stress fluctuations.
@@ II Materials and Methods
The samples are prepared by adding known amount of cetyltrimethyl ammonium tosylate (CTAT) and sodium chloride (NaCl) in filtered deionized water followed by a through mixing inside an air-tight container. After this, the samples are kept in an incubator at (well above the Kraft temperature) for at least 12 hours to ensure complete dissolution of the surfactant in the solution. Further, the solution is equilibrated at room temperature for a week before performing the experiments. As the micellar solution is transparent, to perform particle imaging velocimetry (PIV) measurements, we disperse a known amount (1.5 wt%) of polystyrene tracer particles synthesized in our lab[35] having a size distribution between 20-50 in filtered deionized water before adding CTAT and NaCl. The micellar solution embedded with polystyrene beads is then incubated at for at least 12 hours followed by equilibration at room temperature for a week. The sample is thoroughly mixed before use to ensure a uniform dispersion of the tracer particles. We verify that the addition of 1.5 wt% tracer particles does not change the rheological behavior of the WLM system under study [Fig. S1].
@@
In our experiments, we use a custom-made geometry that is coupled to a MCR-702 Twin-Drive stress-controlled rheometer (Anton-Paar, Graz, Austria ). The geometry has two parts: the measuring geometry and the sample cell. The measuring geometry [Fig. 1(a)-(b)] consists of a millimeter-sized pin (typical length 15-20 mm, thickness/diameter () 1-5 mm) which is connected to the upper drive of the rheometer with the help of a compatible shaft. The sample cell is made from two co-axial cylinders (inner radius () = 58 mm, outer radius () = 83 mm, depth = 18 mm). The pin can be driven inside the sample cell with either force-controlled or velocity-controlled modes (see movie 1) with the help of the rheometer. The samples are filled in the annular region between the cylinders up to a height of 15 mm. We have verified the variation in sample filling height (we have also checked for 10 mm) does not significantly affect the force response of the sample [Fig. S2]. The gap between the end of the cylinder and the bottom plate is approximately 10 mm. Interestingly, we observe that changing the length of the submerged portion of the pin () by varying the gap between the bottom plate and the end of the cylindrical probe does not change the nature of the vs curves. However, increasing the length of the submerged portion causes an upshift of the plots due to the increasing area of contact [Fig. S3]. In our experiments, we have kept the gap between the curved surface of the pin from the side walls (inner and outer walls of the concentric cylinders as shown in Fig.1(a) and 1(b)) also 10 mm. Such an experimental design is compatible both in single drive (only the pin is moving) and in twin drive rheometric configurations (either the pin or the sample cell or both moving). In a single drive mode, the sample cell is mounted on a temperature-controlled Peltier plate. The uncertainty in sample temperature (measured directly using a thermometer) in such configuration is around . The pin is driven inside the sample kept at different temperatures () in order to get the temperature-dependent mechanical response of the viscoelastic fluid. The entire geometry is kept inside a humidity-controlled box to avoid sample evaporation.
@@
For in-situ imaging, we use a twin drive configuration where the sample holder is transparent and can be rotated at a desired velocity with the help of the bottom motor of the rheometer [Fig. S4 and S5]. The pin is held fixed in the lab frame for imaging the fluid around the pin easily. In our set-up direct temperature control of the sample is not possible in such a configuration. We can only control the ambient temperature using an air conditioner (set to ) in the lab. This indirect control increases the uncertainty in sample temperature to around (again confirmed with direct temperature measurement using a thermometer). However, we find that the nature of the mechanical response remains the same within experimental uncertainties. To observe the sample turbidity under imposed flow, we use a transparent sample cell to image the system in the transmission mode in a twin-drive configuration (see Fig. S4 for a schematic of the set-up). As illustrated in the schematic, the illumination is aligned slightly off-axis to the camera position to ensure the collection of only scattered light. To map the local velocity field, we image the sample-air interface (see Fig. S5 for a schematic of the set-up) seeded with tracer beads with the help of a fast camera (Phantom Miro C210). Such images are further analyzed using particle imaging velocimetry algorithm[36] (PIVlab) using MATLAB software to directly map the flow field. PIV analysis is performed by taking a square grid size of 64 pixels and an overlap of 20% between the adjacent grids. The velocity magnitude is calculated from the x and y components followed by a fixed smoothening protocol. There is no further time-averaging to get the instantaneous velocity distribution using PIV.
@@
The extensional behaviour is characterized by DOS (Dripping onto a substrate) measurements where a liquid bridge formed between a dispensing needle and a substrate undergoes uniaxial extensional deformation and self-thinning. For the experiments, we follow standard protocols mentioned in the literature [37, 38]. However, we have replaced the bottom glass substrate with a stainless steel plate of 4 mm diameter to pin the contact line only to the edges of the bottom substrate and avoid continuous spreading of the sample during thinning of the liquid bridge as reported by Omidvar and coworkers [39]. In our experiments, the micellar solution is pumped at a flow rate of 2 with the help of a syringe pump (NE-8000, New Era Pump Systems, USA) and the pumping is stopped before the droplet from the nozzle comes in contact with the substrate. We use a nozzle having an outer diameter of 1.3 mm and an aspect ratio (ratio of the gap between the nozzle and the substrate and nozzle diameter) of 3. The images are recorded at 100 frames per second with the help of a fast camera (Phantom Miro C210) and the data is analyzed using custom written codes in MATLAB software. The experiments are repeated several times to verify the reproducibility of the data.
@@ III Results and discussions
From the raw torque () and angular velocity () signal we obtain the force () and velocity (). Here, corresponds to the pin’s radial distance from the rotation axis as depicted in Fig. 1(b). Fig. 1(c) shows the steady state vs data for 2 wt.% CTAT + 100 mM sodium chloride sample at for pin diameter of 5 mm. We find that varies linearly with up to 10 . At higher velocities, we obtain a plateau region where remains nearly constant with increasing . Beyond this region, we again obtain a linear variation of vs for 100 . The nature of the remains the same for different pins with diameters ranging from 1.5 mm to 5 mm [Fig. S6] except for an upward shift of the curves with increasing pin diameter. In Fig. 1(c) we use 10 s waiting time per data point. We further verify that for larger waiting times , the vs curves remain stationary, indicating the steady state nature of vs data [Fig. S7]. For a comparison we also measure the vs for Newtonian fluid glycerol, where we obtain a linear variation of over a large range of applied velocities [inset of Fig. 1(c)] as expected. We find that for the micellar system is much higher than that for Glycerol for lower pin velocities [Fig. S8]. With increasing velocity beyond the plateau onset, shows a strong decrease due to shear-thinning. Finally, beyond the force-plateau micellar system becomes slightly less than that for glycerol. At such high velocities, the inverse of the shear-rate becomes smaller than the micellar breaking-recombination time scale [16]. This indicates that the broken micellar structures cannot re-join over such a small time scale showing a quasi-Newtonian behaviour (linear vs relation) above the force plateau.
@@
Interestingly, earlier studies on Taylor-Couette flow of CTAT-NaCl systems report the absence of a flat plateau in contrast to the present observation[18, 15, 16]. To test this further, we perform bulk rheological measurements in Taylor-Couette geometry with the CTAT-NaCl WLM system by varying the salt concentrations and temperature. We observe the absence of a flat stress plateau in all cases indicating the non-banding nature of this system (Fig. S9 and S10) over the broad parameter range probed. On the other hand, the presence of a flat force plateau is robustly established with multiple measurements and also by varying the salt concentrations/temperature in the pin-driving setup [Fig. S9, S10 and S11]. Further we calculated the stress due to the pin () by dividing the force over the pin contact area with the sample and shear rate () by dividing the driving velocity over the pin thickness. From the comparison of the flow curves obtained using the pin-driving setup with that observed in bulk rheology, we find a significant mismatch in the nature of the flow curves, as well as, the stress scales [Fig. S12]. This suggests that the nature of the flow and reorganization of WLM around the cylindrical probe is quite different from that of the conventional Taylor-Couette flow and needs more detailed investigations.
@@
Next, we probe the steady-state flow behaviour of the system. We show the temporal evolution of at a constant in Fig. 2(a) for three different imposed velocities: 2.2 mm/s, 9 mm/s, and 15 mm/s. For 15 mm/s which corresponds to the early plateau region of plot, shows large steady-state fluctuations with saw-tooth features, whereas, no significant fluctuations are observed for other values of corresponding to the linear region of vs plot. Interestingly, the saw-tooth nature of fluctuations is observed only at the early plateau region. The nature of the fluctuations becomes more complicated and chaotic deep inside the plateau region [Fig. S13]. Also, the temporal fluctuation of force reaches a steady state fairly quickly across all the driving velocities probed here and do not show any long transients [Fig. S14]. We further quantify the strength of fluctuations by calculating the standard deviation normalized by the mean and find that the strength of fluctuations is maximum exactly at the onset of the plateau region and decreases gradually towards the deeper plateau regime [Fig. S15]. We also confirm that such fluctuations are not observed for Newtonian fluid glycerol over the whole range of that we probe for CTAT-NaCl system [Fig. S16]. This highlights the important role of underlying micellar microstructures and dynamics in generating such unstable flows.

### v2

@@ Abstract
A particle moving through a worm-like micellar fluid (WLM) shows instability and large fluctuations beyond a threshold velocity/force (depending on the control parameter used). Despite many detailed studies, a direct measurement of the time-dependent stress on the probe particle remains unexplored. To address this, we have designed a measuring geometry coupled with a commercial rheometer to study the dynamics of a cylindrical probe through a WLM system of 2 wt.% cetyltrimethyl ammonium tosylate(CTAT) + 100 mM sodium chloride(NaCl) for a wide range of velocity and stress scales. We map out the in-situ velocity distribution using particle imaging velocimetry. Beyond a velocity threshold, we observe large stress fluctuation with gradual stress build-up followed by sudden stress drop indicating storage and release of elastic energy. The length scale constructed from the stress build-up time scale and the probe’s velocity match the length scale of extensile deformation just before the stress drop, confirming the strong correlation of storage and release of energy with the unstable probe motion. Interestingly, the Weissenberg number () for the onset of flow instability obtained from the shear and extensile components remains almost the same. We also find that the turbulent motion of the probe at higher results from the complex mixing of the stick-slip events originating from the partial release of the stored elastic energy. Further, we show that the magnitude of the stick-slip events depends on the detailed micellar structure and dynamics controlled by salt concentration and temperature.
@@ I Introduction
The motion of a particle moving under a constant applied force through a viscous fluid attains a terminal velocity when the viscous drag completely balances the applied force. Such dynamics leading to terminal velocity are well understood[stokes1851effect]. On the other hand, particle motion through visco-elastic worm-like micellar (WLM) fluids can be quite complex, and terminal velocity may not exist due to various fluid instabilities[jayaraman2003oscillations, chen2004flow, kumar2012oscillatory, wu2021linear]. Understanding such complex motion in WLM systems is important not only for a wide range of applications in oil drilling, hydraulic fracturing, and manufacturing cosmetic products, but they also present interesting challenges from the fundamental physics perspectives [bewersdorff1988behaviour, ezrahi2007daily, dreiss2017wormlike, li2020rheological].
@@
Worm-like micelles (WLMs) are long polymer-like structures formed by the self-assembly of surfactant molecules in aqueous solution. In the semi-dilute regime, the density of these polymer-like structures becomes high enough to form entanglement between them. Such entangled structures give rise to long-range correlations in the system resulting in visco-elasticity. Despite many similarities between these systems and polymeric liquids, the major difference between these two classes of systems stems from the fact that WLM can break and recombine due to thermal fluctuations giving rise to additional relaxation modes. This results in Maxwellian behaviour in a wide range of WLM systems with a well-defined relaxation time scale [cates1987reptation]. Non-linear flows of WLM systems have been extensively studied using shear rheometry as they exhibit striking flow instabilities, such as shear banding, elastic turbulence, rheochaos at substantially low values of Reynolds numbers but at high Weissenberg numbers [salmon2003velocity, helgeson2009relating, mohammadigoushki2016flow, fardin2010elastic, majumdar2011universality, ghadai2023origin, bandyopadhyay2000observation, ganapathy2006intermittency, ganapathy2008spatiotemporal]. Such instabilities are driven by elasticity rather than inertia. Nonetheless, interesting inertio-elastic flow instabilities have also been reported in these systems when both and becomes comparable [perge2014inertio, mohammadigoushki2017inertio]. Although these studies provide valuable insights regarding the coupling of microstructure and flow in WLM systems, such rheological studies are inadequate to predict/describe the complex motion of a probe through WLM fluids mostly due to the wide differences in flow geometries.
@@
Earlier experimental studies have reported the unstable motion of a sphere settling under gravity through WLM fluids. These include the study by Jayraman and Belmonte[jayaraman2003oscillations] using CTAB-NaSal based WLM fluid. They observe that beyond a certain value of Deborah number 45, the velocity of the sphere shows large fluctuations. They have attributed the origin of such fluctuations to the formation and breakage of shear-induced structures around the sphere. More recent studies by Mohammadigoushki and Muller[mohammadigoushki2016sedimentation] and Zhang and Muller[zhang2018unsteady] point out the role of extensional deformation in the wake of the moving sphere giving rise to instability beyond the certain value of the extensional Weissenberg number. Using in-situ flow birefringence studies Chen and Rothstein[chen2004flow] suggest that the flow instability originates from the flow-induced rupture of stretched micelles at the rear end of the moving sphere. They also directly observe filament rupture using extensional rheometry. The experimental investigation of Wu and Mohammadigoushki[wu2021linear] further supports a similar mechanism of micellar breakage. Kumar et al.[kumar2012oscillatory] show unstable flows in CTAT + NaCl based WLM system where they report a periodic burst of damped oscillations in the sphere’s velocity. The nature of such oscillations is quite distinct from the similar studies mentioned above. This implies that the microscopic micellar structure plays an important role in determining the nature of instability. Nonetheless, the steep rise in velocity of the settling sphere at the onset of a burst also suggests a sudden structural breakage. Interestingly, a recent bulk rheological study on the CTAT-NaCl WLM system highlights the role of micellar breakage in the elasticity-driven Taylor-Couette instability in these systems[ghadai2023origin]. Recent numerical simulation based on two-species VCM model[vasquez2007network] for WLM also supports the micellar rupture in the wake of a moving sphere as the mechanism of instability[sasmal2021unsteady]. The role of surface roughness of the settling sphere has also been investigated in WLM systems[sasmal2022effect]. The structure of CTAT-based WLM systems is complex and highly sensitive to the nature and concentration of added salt which in turn affect their rheological response significantly[calabrese2015rheology, rassolov2022kinetics]. Very recent studies indicate that WLM systems having almost identical linear rheological properties can display significantly different non-linear dynamical behaviour[pasquino2023startup].
@@
Despite these detailed studies, a direct measurement of temporal stress on the probe during the transition from stable to unstable motion by systematically varying the flow rate remains unexplored. Based on pressure drop measurements in an array of cylinders [moss2010flow] and position fluctuations of a flexible elastomer [dey2018viscoelastic] placed in a flow-field of WLM fluids, the role of stress variation during the flow instabilities have been indicated. Nonetheless, stress measurement with high spatial and temporal resolution is crucial to verify the proposed mechanisms of micellar scission in giving rise to unstable motion through WLM systems. Furthermore, such controlled experiments are required to understand the connection of the instability near the onset point to the chaotic or turbulent motion at much higher values of as reported in the earlier studies for CTAT-NaCl systems.
@@
In this work, we use a home-built setup coupled with a commercial rheometer to study the stress response of a cylindrical probe driven through the WLM system with constant velocity. Our setup not only mimics the motion of small probes settling under gravity in a more controlled manner but also has other significant advantages: For a given WLM system, the wide range of velocity/force scales spanning several decades that can be probed/measured in our pin-driving setup can not be matched in conventional ball-drop experiments. Similarly, in our setup the stress fluctuation data can be acquired, as long as the sample remains stable. This allows us to perform accurate statistical analysis over very long time scales. Thus, our setup can potentially verify the applicability of the existing macroscopic rheological models of flow instability, transient dynamics, shear banding etc.[mohammadigoushki2019transient, perge2014surfactant, divoux2016shear] in the context of unstable motion of a small probe through similar systems. This setup also allows us to probe the in-situ local deformation field using PIV and light scattering measurements. We observe the formation of a turbid band in the wake of the probe, just below the onset velocity of the probe corresponding to large stress fluctuations. We find that such a turbid band corresponds to an extended wake and its attachment-detachment dynamics with the probe is strongly correlated with the nature of the stress fluctuations.
@@ II Materials and Methods
The samples are prepared by adding known amount of cetyltrimethyl ammonium tosylate (CTAT) and sodium chloride (NaCl) in filtered deionized water followed by a through mixing inside an air-tight container. After this, the samples are kept in an incubator at (well above the Kraft temperature) for at least 12 hours to ensure complete dissolution of the surfactant in the solution. Further, the solution is equilibrated at room temperature for a week before performing the experiments. As the micellar solution is transparent, to perform particle imaging velocimetry (PIV) measurements, we disperse a known amount (1.5 wt%) of polystyrene tracer particles synthesized in our lab[dhar2019signature] having a size distribution between 20-50 in filtered deionized water before adding CTAT and NaCl. The micellar solution embedded with polystyrene beads is then incubated at for at least 12 hours followed by equilibration at room temperature for a week. The sample is thoroughly mixed before use to ensure a uniform dispersion of the tracer particles. We verify that the addition of 1.5 wt% tracer particles does not change the rheological behavior of the WLM system under study [Fig. S1].
@@
In our experiments, we use a custom-made geometry that is coupled to a MCR-702 Twin-Drive stress-controlled rheometer (Anton-Paar, Graz, Austria ). The geometry has two parts: the measuring geometry and the sample cell. The measuring geometry [Fig. 1(a)-(b)] consists of a millimeter-sized pin (typical length 15-20 mm, thickness/diameter () 1-5 mm) which is connected to the upper drive of the rheometer with the help of a compatible shaft. The sample cell is made from two co-axial cylinders (inner radius () = 58 mm, outer radius () = 83 mm, depth = 18 mm). The pin can be driven inside the sample cell with either force-controlled or velocity-controlled modes (see movie 1) with the help of the rheometer. The samples are filled in the annular region between the cylinders up to a height of 15 mm. We have verified the variation in sample filling height (we have also checked for 10 mm) does not significantly affect the force response of the sample [Fig. S2]. The gap between the end of the cylinder and the bottom plate is approximately 10 mm. Interestingly, we observe that changing the length of the submerged portion of the pin by varying the gap between the bottom plate and the end of the cylindrical probe does not change the nature of the vs curves. However, increasing the length of the submerged portion causes an upshift of the plots due to the increasing area of contact [Fig. S3]. In our experiments, we have kept the gap between the curved surface of the pin from the side walls (inner and outer walls of the concentric cylinders as shown in Fig.1(a) and 1(b)) also 10 mm. Such an experimental design is compatible both in single drive (only the pin is moving) and in twin drive rheometric configurations (either the pin or the sample cell or both moving). In a single drive mode, the sample cell is mounted on a temperature-controlled Peltier plate. The uncertainty in sample temperature (measured directly using a thermometer) in such configuration is around . The pin is driven inside the sample kept at different temperatures () in order to get the temperature-dependent mechanical response of the viscoelastic fluid. The entire geometry is kept inside a humidity-controlled box to avoid sample evaporation.
@@
For in-situ imaging, we use a twin drive configuration where the sample holder is transparent and can be rotated at a desired velocity with the help of the bottom motor of the rheometer [Fig. S4 and S5]. The pin is held fixed in the lab frame for imaging the fluid around the pin easily. In our set-up direct temperature control of the sample is not possible in such a configuration. We can only control the ambient temperature using an air conditioner (set to ) in the lab. This indirect control increases the uncertainty in sample temperature to around (again confirmed with direct temperature measurement using a thermometer). However, we find that the nature of the mechanical response remains the same within experimental uncertainties. To observe the sample turbidity under imposed flow, we use a transparent sample cell to image the system in the transmission mode in a twin-drive configuration (see Fig. S4 for a schematic of the set-up). As illustrated in the schematic, the illumination is aligned slightly off-axis to the camera position to ensure the collection of only scattered light. To map the local velocity field, we image the sample-air interface (see Fig. S5 for a schematic of the set-up) seeded with tracer beads with the help of a fast camera (Phantom Miro C210). Such images are further analyzed using particle imaging velocimetry algorithm[Thielicke_2021] (PIVlab) using MATLAB software to directly map the flow field. PIV analysis is performed by taking a square grid size of 64 pixels and an overlap of 20% between the adjacent grids. The velocity magnitude is calculated from the x and y components followed by a fixed smoothening protocol. There is no further time-averaging to get the instantaneous velocity distribution using PIV.
@@
The extensional behaviour is characterized by DOS (Dripping onto a substrate) measurements where a liquid bridge formed between a dispensing needle and a substrate undergoes uniaxial extensional deformation and self-thinning. For the experiments, we follow standard protocols mentioned in the literature [dinic2015extensional, dinic2017pinch]. However, we have replaced the bottom glass substrate with a stainless steel plate of 4 mm diameter to pin the contact line only to the edges of the bottom substrate and avoid continuous spreading of the sample during thinning of the liquid bridge as reported by Omidvar and coworkers [omidvar2019detecting]. In our experiments, the micellar solution is pumped at a flow rate of 2 with the help of a syringe pump (NE-8000, New Era Pump Systems, USA) and the pumping is stopped before the droplet from the nozzle comes in contact with the substrate. We use a nozzle having an outer diameter of 1.3 mm and an aspect ratio (ratio of the gap between the nozzle and the substrate and nozzle diameter) of 3. The images are recorded at 100 frames per second with the help of a fast camera (Phantom Miro C210) and the data is analyzed using custom written codes in MATLAB software. The experiments are repeated several times to verify the reproducibility of the data.
@@ III Results and discussions
From the raw torque () and angular velocity () signal we obtain the force () and velocity (). Here, corresponds to the pin’s radial distance from the rotation axis as depicted in Fig. 1(b). Fig. 1(c) shows the steady state vs data for 2 wt.% CTAT + 100 mM sodium chloride sample at for pin diameter of 5 mm. We find that varies linearly with up to 10 . At higher velocities, we obtain a plateau region where remains nearly constant with increasing . Beyond this region, we again obtain a linear variation of vs for 100 . The nature of the remains the same for different pins with diameters ranging from 1.5 mm to 5 mm [Fig. S6] except for an upward shift of the curves with increasing pin diameter. In Fig. 1(c) we use 10 s waiting time per data point. We further verify that for larger waiting times , the vs curves remain stationary, indicating the steady state nature of vs data [Fig. S7]. However, there is a slight deviation of at the plateau onset over repeated measurements. This could happen due to the presence of large fluctuations as discussed in the subsequent sections of the paper. We also measure the vs for Newtonian fluid glycerol, where we obtain a linear variation of over a large range of applied velocities [inset of Fig. 1(c)] as expected. We find that for the micellar system is much higher than that for Glycerol for lower pin velocities [Fig. S8]. With increasing velocity beyond the plateau onset, shows a strong decrease due to shear-thinning. Finally, beyond the force-plateau micellar system becomes slightly less than that for glycerol. At such high velocities, the inverse of the shear-rate becomes smaller than the micellar breaking-recombination time scale [ghadai2023origin]. This indicates that the broken micellar structures cannot re-join over such a small time scale showing a quasi-Newtonian behaviour (linear vs relation) above the force plateau.
Interestingly, earlier studies on Taylor-Couette flow of CTAT-NaCl systems report the absence of a flat stress plateau [ganapathy2006intermittency, majumdar2011universality, ghadai2023origin]. The significant power-law slope ( 0.3) at the stress plateau indicates that the CTAT-NaCl system is not expected to exhibit shear banding. We perform bulk rheological measurements in Taylor-Couette geometry with the CTAT-NaCl WLM system by varying the salt concentrations and
temperature. We observe the absence of a flat stress plateau
in all cases (Fig. S9 and S10) over the broad parameter range probed further supporting the non-banding nature of the system. The evidence of stable shear bands is also not observed during flow-induced turbidity measurements in the gradient-vorticity plane [ghadai2023origin]. However, from the PIV studies in the flow-gradient plane, we find that in the early plateau region, before
the onset of elastic turbulence, the velocity profiles across the shear gap show a time variation with the velocity fluctuations predominantly confined close to the rotating inner cylinder [ghadai2023origin]. This observation suggests that the possibility of flow-induced
@@
spatial heterogeneities in the system cannot be ruled out, although no steady shear bands are observed. On the other hand, the presence of a flat force plateau is robustly established with multiple measurements and also by varying the salt concentrations/temperature in the pin-driving setup [Fig. S9, S10 and S11]. Further, we calculated the stress due to the pin () by dividing the force over the pin contact area with the sample and shear rate () by dividing the driving velocity over the pin thickness. From the comparison of the flow curves obtained using the pin-driving setup with that observed in bulk rheology, we find a significant mismatch in the nature of the flow curves, as well as, the stress scales [Fig. S12]. This suggests that the nature of the flow and reorganization of WLM around the cylindrical probe is quite different from that of the conventional Taylor-Couette flow and needs more detailed investigations.
@@
Next, we probe the steady-state flow behaviour of the system. We show the temporal evolution of at a constant in Fig. 2(a) for three different imposed velocities: 2.2 mm/s, 9 mm/s, and 15 mm/s. For 15 mm/s which corresponds to the early plateau region of plot, shows large steady-state fluctuations with saw-tooth features, whereas, no significant fluctuations are observed for other values of corresponding to the linear region of vs plot. Interestingly, the saw-tooth nature of fluctuations is observed only at the early plateau region. The nature of the fluctuations becomes more complicated and chaotic deep inside the plateau region [Fig. S13]. Also, the temporal fluctuation of force reaches a steady state fairly quickly across all the driving velocities probed here and does not show any long transients [Fig. S14]. To rule out long transients in the time series and viscous heating, we divide the force time series (obtained for at least 2000 seconds for each driving velocity) in to 4 equal parts (500 seconds in each part) and plot the mean force with error bars indicating the standard deviations [Fig. S15]. We find that the mean and the standard deviation remain the same across the different partitions of the time series. This indicates statistical stationarity, implying a steady state behaviour. We further quantify the strength of fluctuations by calculating the standard deviation normalized by the mean and find that the strength of fluctuations is maximum exactly at the onset of the plateau region and decreases gradually towards the deeper plateau regime [Fig. S16]. We also confirm that such fluctuations are not observed for Newtonian fluid glycerol over the whole range of that we probe for CTAT-NaCl system [Fig. S17]. This highlights the important role of underlying micellar microstructures and dynamics in generating such unstable flows.